OpenStudy (anonymous):

help 0. Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

what willl be slope?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

first take equ in standard form y=mx+c now slope?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

no 2x-3y=12 -3y=-2x+12 y=2/3x-4 slope ?

OpenStudy (anonymous):

3x-4 or 2?

OpenStudy (anonymous):

2/3 is coefficient of x so slope =2/3

OpenStudy (anonymous):

slope of perpendicular is negative invese of this slope I taught u earlier so slope of per line

OpenStudy (anonymous):

ok can u help me with number 1 n 2? imma attatch the file bc it wont let me attactch the pic so here ITS # 1 N 2

OpenStudy (anonymous):

slope =y/x =-2/-5=2/5 equ y=mx+b b is y intercept b=-2 put m and b in equ and solve

OpenStudy (anonymous):

wait wat does m =?

OpenStudy (anonymous):

for 2 eq is x=3

OpenStudy (anonymous):

m slope 2/5

OpenStudy (anonymous):

its hard :(

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

@hartnn can u help me?

OpenStudy (anonymous):

ok can u help me with number 1 n 2? imma attatch the file bc it wont let me attactch the pic so here ITS # 1 N 2 7 days ago 1 attachments writingequations_writingassignment1.pdf

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Kira_Yamato can u help me with numbers 1 and 2 in the attached file?

OpenStudy (kira_yamato):

1) y-intercept is -2 Gradient = (0 - (-2)) / (-6-0) = 2/(-6) = -1/3 2) x = 3 (since vertical line intersecting x-axis at 3)